# Robust Cross-View Geo-Localization via Content-Viewpoint Disentanglement

### v1

@@ 3.3 Intra-view Independence Constraint
where $\mathcal{SW}_{2}(\cdot,\cdot)$ denotes the Sliced Wasserstein-2 distance. We apply this constraint independently to both views, yielding $\mathcal{L}_{\text{iic}}^{\mathrm{q}}$ and $\mathcal{L}_{\text{iic}}^{\mathrm{s}}$. This constraint drives the separation of view-agnostic scene structure from view-specific attributes in a computationally tractable manner.
@@ 3.4 Inter-view Reconstruction Constraint
While the independence constraint promotes factor separation, it does not guarantee that each embedding retains the essential factor-specific information. In particular, relying solely on independence may lead to trivial solutions where either the content or viewpoint embedding becomes uninformative.
@@
^$\hat{\mathcal{I}}^{\mathrm{q}}=D^{\mathrm{q}}(\mathbf{z}_{c}^{\mathrm{s}},\mathbf{z}_{v}^{\mathrm{q}}),\quad\hat{\mathcal{I}}^{\mathrm{s}}=D^{\mathrm{s}}(\mathbf{z}_{c}^{\mathrm{q}},\mathbf{z}_{v}^{\mathrm{s}}).$ (4)
@@
where $\|\cdot\|_{2}^{2}$ denotes the mean squared error (MSE) between the original and reconstructed images. This loss is also applied to both views, resulting in $\mathcal{L}_{\text{irc}}^{\mathrm{q}}$ and $\mathcal{L}_{\text{irc}}^{\mathrm{s}}$, which ensure that the disentangled features preserve the necessary information to reconstruct their cross-view counterparts.
@@ 3.5 Training Objective
Following prior works [4, 11], we employ the standard InfoNCE loss for view-agnostic content consistency across views, denoted as $\mathcal{L}_{\text{loc}}$, which encourages content of matched query-satellite pairs to be close while pushing away mismatched pairs, and is defined as:
$\mathcal{L}_{\text{loc}}=-\log\frac{\exp(\mathbf{z}_{c}^{\mathrm{q}}\cdot\mathbf{z}_{c}^{\mathrm{s}}/\tau)}{\sum_{i=1}^{N}\exp(\mathbf{z}_{c}^{\mathrm{q}}\cdot\mathbf{z}_{c}^{\mathrm{i}}/\tau)},$ (6)
@@
$\mathcal{L}_{\text{total}}=\lambda_{1}(\tfrac{1}{2}\mathcal{L}^{\mathrm{q}}_{\text{iic}}+\tfrac{1}{2}\mathcal{L}^{\mathrm{s}}_{\text{iic}})+\lambda_{2}(\tfrac{1}{2}\mathcal{L}^{\mathrm{q}}_{\text{irc}}+\tfrac{1}{2}\mathcal{L}^{\mathrm{s}}_{\text{irc}})+\lambda_{3}\mathcal{L}_{\text{loc}},$ (7)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters that balance the three losses during training.
@@ 4.1 Settings
Datasets. We evaluate our method on four representative CVGL benchmarks: University-1652 [28], SUES-200 [30], CVUSA [21], and CVACT_val [9], which collectively encompass various perspective combinations, including ground-view, drone-view, and satellite-view imagery. University-1652 comprises images from 1,652 university campuses captured in ground, drone, and satellite views. In our experiments, we adopt the drone-satellite setting, using 701 campuses for training, 701 for testing, along with 250 distractor samples. SUES-200 consists of real-world drone and satellite imagery from 200 scenes across four altitudes (150-300m), with 120 scenes for training and 80 for testing. CVUSA and CVACT_val provide large-scale ground-satellite image pairs, with 35,532 aligned in the training set for each. The former contains 8,884 test queries, while the latter offers a validation set of the same size.
Evaluation metrics. We adopt five standard retrieval metrics, including Average Precision (AP), Recall@K (K=1,5,10), and Recall@1%, to evaluate cross-view matching performance. Detailed definitions of metrics can be found in the Appendix.
Implementation details. The hyperparameters of the total training loss $\mathcal{L}_{\text{total}}$ are fixed across all models and datasets, with $\lambda_{1}=10$, $\lambda_{2}=0.2$, and $\lambda_{3}=1$. The temperature parameter $\tau$ is set to 0.05. To ensure fair comparisons, we strictly follow the original training configurations of each baseline, including optimizer type, learning rate schedule, batch size, etc., without any additional tuning. All experiments are conducted in PyTorch on an NVIDIA RTX 4090 GPU with 24 GB memory.
4.2 Results on University-1652
[…]
Table 1: Comparison between CVD and baseline methods on University-1652. $\dagger$’ represents that CVD is applied during training.
@@ 4.2 Results on University-1652
MCCG 256$\times$256 ConvNeXt-T 90.63 88.92 96.44 97.63 97.77 88.73 93.15 95.72 96.72 99.57
@@
Sample4Geo 384$\times$384 ConvNeXt-B 93.56 92.36 97.64 98.26 98.36 91.64 94.72 97.00 97.43 99.43
@@
Table 2: Comparison between CVD and baseline methods on SUES-200. $\dagger$’ represents that CVD is applied during training.
@@ 4.3 Results on SUES-200
4.4 Results on CVUSA and CVACT
We further evaluate CVD on CVUSA and CVACT, which involve ground-to-satellite matching under extreme viewpoint differences and substantial scene layout variation. As shown in Table 3, all baselines exhibit measurable improvements when trained with CVD, even though they already perform strongly. These improvements are particularly meaningful given the challenges posed by ground-view imagery, such as occlusion, illumination changes, and perspective distortion.
Table 3: Comparison between CVD and baseline methods on CVUSA and CVACT. $\dagger$’ represents that CVD is applied during training.
@@ 4.4 Results on CVUSA and CVACT
Sample4Geo $\dagger$ 98.67 99.40 99.78 99.89 90.94 96.80 97.51 98.81
By explicitly separating view-specific conflicting information, CVD enhances the consistency of content representations across modalities. For example, Sample4Geo improves by +0.24% and +0.65% in R@1 on CVUSA and CVACT, respectively, while GeoDTR achieves +0.77% and +0.61% gains in R@10.
Figure 3: Visualization of disentangled content and viewpoint attention maps produced by CVD.
4.5 Visualization of Disentangled Embeddings
Figure 3 visualizes attention maps derived from the disentangled embeddings produced by CVD. Viewpoint embeddings predominantly focus on peripheral, view-specific elements such as vegetation, shadows, and roads, which are often inconsistent across modalities. In contrast, content embeddings attend to semantically stable regions like building structures, providing reliable and discriminative cues for cross-view matching. This separation facilitates CVD to learn robust, viewpoint-agnostic content representations, thereby improving retrieval under significant viewpoint disparity. More visualizations are given in the Appendix.
Table 4: Ablation of the components of CVD.
Exp. Methods (ResNet50) D $\rightarrow$ S S $\rightarrow$ D
@@ 4.5 Visualization of Disentangled Embeddings
3 Exp.1 + iic loss 77.95 73.95 73.97 85.24
4 Exp.1 + irc loss 78.28 74.62 74.63 85.45
@@
8 Exp.6 + iic loss 73.29 69.19 66.70 83.20
9 Exp.7 + irc loss 73.60 69.78 67.03 83.61
@@
Table 5: Ablation of content-viewpoint ratio.
Split Ratio D $\rightarrow$ S S $\rightarrow$ D
@@
Table 6: Influence of different $\tau$ in InfoNCE.
Method D $\rightarrow$ S S $\rightarrow$ D
@@ 4.6 Ablation Study
Effectiveness of CVD. We conduct ablation studies on University-1652 to evaluate the contribution of each component in CVD, as summarized in Section 4.5. To ensure fair comparison, we adopt two representative pipelines, LPN and Game4Loc, which both use a shared ResNet50 backbone. Removing both constraints (Exp.2 and 7) leads to a notable performance drop, while individually adding the intra-view independence (Exp.3 and 8) or inter-view reconstruction constraint (Exp.4 and 9) yields consistent gains. The best results are obtained when both constraints are jointly applied (Exp.5 and 10), confirming that explicitly factorizing content and viewpoint is essential for robust cross-view alignment.
@@
Temperature Parameter Analysis. As shown in Section 4.5, model performance remains stable across a range of temperature values $\tau$. This insensitivity indicates that the performance gains primarily result from effective representation disentanglement rather than contrastive loss tuning.
Table 7: Generalization ablation when trained on University-1652 and tested on SUES-200. $\dagger$’ represents that CVD is applied during training.
Method Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
150m 200m 250m 300m 150m 200m 250m 300m
AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1
Game4Loc $\dagger$ 86.12 82.87 91.77 89.70 93.59 91.92 94.44 92.92 75.37 87.50 84.25 90.00 87.97 95.00 90.21 95.00
4.7 Generalisation Capabilities
To evaluate the generalization ability of CVD to unseen scenes, we train models on University-1652 and directly evaluate them on SUES-200 without any fine-tuning. As shown in Table 7, CVD significantly improves performance across all metrics for both LPN (ResNet-50) and Game4Loc (ViT-B). For example, it boosts R@1 by up to +6.25% for LPN and +7.5% for Game4Loc, even without access to the target domain during training. Remarkably, these cross-dataset improvements exceed those typically obtained via in-domain training on SUES-200, which yields only improvements of 1-3%. These results highlight that disentangling content and viewpoint leads to more transferable representations that generalize effectively across regions and views.
Figure 4: Cross-view reconstructions with matched (left column) and unrelated (right column) viewpoints.
4.8 Effectiveness of Content-Viewpoint Disentanglement
To further evaluate the effectiveness of disentangling content and viewpoint, we conduct a cross-view recombination experiment using the learned embeddings. As illustrated in Figure 4, each row shows two reconstructed heatmaps, which share the same content but differ in the viewpoint. In the top row, the same satellite content embedding is combined with either a matched (left) or unrelated (right) drone viewpoint; in the bottom row, drone content is paired with matched and unrelated satellite viewpoints, respectively. Despite the change in viewpoint, the resulting attention maps exhibit similar attention patterns, highlight semantically consistent regions. These results provide qualitative evidence that CVD effectively disentangles the two factors: content embeddings encode view-agnostic structural information, while viewpoint embeddings modulate geometric and contextual variations. More visualizations are provided in the Appendix.
5 Conclusion and Limitations
In this paper, we revisit cross-view geo-localization (CVGL) from a manifold learning perspective and propose CVD, a unified framework that explicitly disentangles content and viewpoint within the visual representation. CVD adopts an embed-disentangle-reconstruct paradigm, guided by intra-view independence and inter-view reconstruction constraints, to promote content- and viewpoint-specific encoding. Extensive experiments demonstrate that CVD consistently enhances both localization accuracy and generalization across diverse CVGL pipelines. These findings underscore the value of explicitly separating content and viewpoint in CVGL, and offer insights toward developing more robust and generalizable CVGL systems. A discussion of limitations and potential future directions is provided in the Appendix.

### v2

@@ 3.3 Intra-view Independence Constraint
where $\mathcal{SW}_{2}(\cdot,\cdot)$ denotes the Sliced Wasserstein-2 distance. We apply this constraint independently to both views, yielding $\mathcal{L}_{\text{iic}}^{\mathrm{d}}$ and $\mathcal{L}_{\text{iic}}^{\mathrm{s}}$. This constraint drives the separation of view-agnostic scene structure from view-specific attributes in a computationally tractable manner.
@@ 3.4 Inter-view Reconstruction Constraint
While the independence constraint promotes factor separation, it does not guarantee that each embedding retains the essential factor-specific information. In particular, relying solely on independence may lead to trivial solutions where either the content or the viewpoint embedding becomes uninformative.
@@
$\hat{\mathcal{I}}^{\mathrm{d}}=D^{\mathrm{d}}(\mathbf{z}_{c}^{\mathrm{s}},\mathbf{z}_{v}^{\mathrm{d}}),\quad\hat{\mathcal{I}}^{\mathrm{s}}=D^{\mathrm{s}}(\mathbf{z}_{c}^{\mathrm{d}},\mathbf{z}_{v}^{\mathrm{s}}).$ (4)
@@
where $\|\cdot\|_{2}^{2}$ denotes the mean squared error (MSE) between the original and reconstructed images. This loss is also applied to both views, resulting in $\mathcal{L}_{\text{irc}}^{\mathrm{d}}$ and $\mathcal{L}_{\text{irc}}^{\mathrm{s}}$, which ensure that the disentangled features preserve the necessary information to reconstruct their cross-view counterparts.
@@ 3.5 Training Objective
Following prior works [deuser2023sample4geo, mi2024congeo], we employ the standard InfoNCE loss for view-agnostic content consistency across views, denoted as $\mathcal{L}_{\text{loc}}$, which encourages content of matched drone-satellite pairs to be close while pushing away mismatched pairs, and is defined as:
$\mathcal{L}_{\text{loc}}=-\log\frac{\exp(\mathbf{z}_{c}^{\mathrm{d}}\cdot\mathbf{z}_{c}^{\mathrm{s}}/\tau)}{\sum_{i=1}^{N}\exp(\mathbf{z}_{c}^{\mathrm{d}}\cdot\mathbf{z}_{c}^{\mathrm{i}}/\tau)},$ (6)
@@
$\mathcal{L}_{\text{total}}=\lambda_{1}(\tfrac{1}{2}\mathcal{L}^{\mathrm{d}}_{\text{iic}}+\tfrac{1}{2}\mathcal{L}^{\mathrm{s}}_{\text{iic}})+\lambda_{2}(\tfrac{1}{2}\mathcal{L}^{\mathrm{d}}_{\text{irc}}+\tfrac{1}{2}\mathcal{L}^{\mathrm{s}}_{\text{irc}})+\mathcal{L}_{\text{loc}},$ (7)
where $\lambda_{1}$ and $\lambda_{2}$ are two loss-balancing hyperparameters.
@@ 4.1 Settings
Datasets. We evaluate our method on four representative CVGL benchmarks: University-1652 [zheng2020university] comprises images from 1,652 university campuses captured in ground, drone, and satellite views. In our experiments, we adopt the drone-satellite setting, using 701 campuses for training, 701 for testing, along with 250 distractor samples. SUES-200 [zhu2023sues] consists of real-world drone and satellite imagery from 200 scenes across four altitudes (150-300m), with 120 scenes for training and 80 for testing. Together, University-1652 and SUES-200 span diverse scene types, flight altitudes, and viewing directions, providing a rigorous testbed to assess the robustness and generalization of our approach. To further validate effectiveness, we additionally evaluate on the CVUSA [workman2015wide] and CVACT_val [liu2019lending] datasets (ground $\rightarrow$ satellite), which provide 35,532 aligned training pairs each; the former contains 8,884 test queries, and the latter offers a test set of the same size.
Evaluation metrics. We adopt five retrieval metrics, including Average Precision (AP), Recall@K (K=1,5,10), and Recall@1%, to evaluate cross-view matching performance. Definitions of these metrics are provided in the Appendix.
Baselines. To evaluate the effectiveness of our method, we compare it against several SOTA methods, including LPN [wang2021each], FSRA [dai2021transformer], TransGeo [zhu2022transgeo], MCCG [shen2023mccg], Sample4Geo [deuser2023sample4geo], SDPL [chen2024sdpl], Game4Loc [ji2025game4loc] and GeoDTR [zhang2023cross].
Implementation details. The hyperparameters of the total training loss are fixed across all experiments, with $\lambda_{1}=10$, and $\lambda_{2}=0.2$. The temperature parameter $\tau$ is set to 0.05. To ensure fair comparisons, we strictly follow the original training configurations of each baseline, including optimizer type, learning rate schedule, etc., without any additional tuning. All experiments are conducted in PyTorch on an NVIDIA 4090 GPU. Each experiment is repeated three times with different random seeds, and the mean results are reported to ensure statistical reliability.
Table 1: Comparison of baselines and their CVD-enhanced counterparts (marked with $\dagger$) on the University-1652 dataset.
@@
SDPL $\times512$ ResNet50 86.45 84.13 94.36 96.45 96.72 82.17 89.44 92.58 93.58 99.29
@@
MCCG $\times256$ ConvNeXt-T 90.63 88.92 96.44 97.63 97.77 88.73 93.15 95.72 96.72 99.57
@@
Game4Loc $\times384$ ViT-B 92.56 91.32 96.56 97.33 97.41 90.83 94.43 95.72 96.57 98.71
[…]
Table 2: Comparison of baselines and their CVD-enhanced counterparts (marked with $\dagger$) on the SUES-200 dataset.
@@
4.2 Main Results
[…]
Table 3: Cross-dataset generalization results comparing baselines and their CVD-enhanced counterparts (marked with $\dagger$), trained on the University-1652 dataset and directly tested on the SUES-200 dataset.
Method Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
150m 200m 250m 300m 150m 200m 250m 300m
AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1
FSRA 58.22 52.45 67.10 61.87 70.64 66.07 71.99 67.50 50.95 58.75 59.07 66.25 61.07 62.50 61.98 63.75
[…]
SDPL 38.52 32.80 47.30 41.33 52.38 46.72 53.62 48.50 25.26 27.50 35.05 37.50 41.74 46.25 43.85 48.75
[…]
MCCG 74.99 70.85 86.04 83.20 90.84 88.90 93.38 91.85 59.85 63.75 74.65 81.25 79.87 83.75 81.17 86.25
[…]
Sample4Geo 64.24 57.62 74.45 69.00 81.22 77.02 85.48 81.90 83.85 88.75 90.09 92.50 91.68 96.25 93.51 95.00
[…]
Table 4: Comparison of baselines and their CVD-enhanced counterparts (marked with $\dagger$) on the CVUSA and CVACT datasets.
@@ 4.2 Main Results
[…]
Results on University-1652 $\rightarrow$ SUES-200. To evaluate the generalization ability of CVD to unseen scenes, we train models on University-1652 and directly evaluate them on SUES-200 without any fine-tuning. As shown in Tab. 3, CVD significantly improves performance across all metrics for both LPN (ResNet-50) and Game4Loc (ViT-B). For example, it boosts R@1 by up to +6.25% for LPN and +7.5% for Game4Loc, even without access to the target domain during training. Remarkably, these cross-dataset improvements exceed those typically obtained via in-domain training on SUES-200, which yields only improvements of 1-3%. These results highlight that disentangling content and viewpoint leads to more transferable representations that generalize effectively across regions and views.
Results on CVUSA and CVACT. To further examine the generality of CVD beyond the drone-view setting, we also evaluate it on CVUSA and CVACT, which involve ground-to-satellite matching under extreme viewpoint differences and substantial scene layout variation. As shown in Tab. 4, all baselines exhibit continuous improvements when trained with CVD, even though they already perform strongly. These improvements are particularly meaningful given the challenges posed by ground-view imagery, such as occlusion, illumination changes, and perspective distortion. By explicitly separating view-specific conflicting information, CVD enhances the consistency of content representations across modalities. For example, Sample4Geo improves by +0.24% and +0.65% in R@1 on CVUSA and CVACT, respectively, while GeoDTR achieves +0.77% and +0.61% gains in R@10.
Qualitative Results. Appendix Sec. 10.1, we provide qualitative comparisons of cross-view retrieval results on University-1652 using both CNN-based LPN and Transformer-based Game4Loc. The results intuitively confirm that CVD disentangles view-agnostic content from viewpoint-specific variations, enabling more reliable cross-view correspondence.
[…]
Figure 3: Qualitative results of cross-view reconstruction on the University-1652 and SUES-200 datasets.
Table 5: Ablation studies on: (a) Effect of CVD’s components. (b) Content-viewpoint split ratio $\alpha$. (c) Analysis of different $\tau$ in InfoNCE.
Exp. Methods (ResNet50) Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
@@
3 Exp.1 + $\mathcal{L}_{\text{iic}}$ 77.95 73.95 73.97 85.24
4 Exp.1 + $\mathcal{L}_{\text{irc}}$ 78.28 74.62 74.63 85.45
@@
8 Exp.6 + $\mathcal{L}_{\text{iic}}$ 73.29 69.19 66.70 83.20
9 Exp.7 + $\mathcal{L}_{\text{irc}}$ 73.60 69.78 67.03 83.61
@@
Split Ratio Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
@@
Method Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
@@
Table 6: Effect of different reconstruction loss functions on the University-1652 dataset.
[…]
AP R@1 R@5 R@10 R@1% AP R@1 R@5 R@10 R@1%
MSE 19.38 0.4608 93.95 92.94 97.59 98.28 98.32 91.92 94.86 96.71 97.14 99.57
SSIM 19.21 0.4601 93.19 92.05 97.26 97.87 97.90 91.78 94.55 96.40 96.19 99.56
Perceptual 19.06 0.4589 92.70 91.45 96.82 97.84 97.91 90.85 94.14 97.42 97.26 99.29
4.3 Effectiveness of Disentangled Strategy
To evaluate our disentanglement strategy, we conduct cross-view image reconstruction (Drone $\rightarrow$ Satellite) on University-1652 and SUES-200. As illustrated in Fig. 3, the visualization results show that the reconstructed outputs consistently preserve the global layout, semantic topology, and structural relations of the original scenes, despite the loss of certain high-frequency and color details. For example, the upper-right sample faithfully recovers the circular central plaza and the relative arrangement of surrounding buildings. This indicates that CVD successfully learned meaningful content and viewpoint representations, thereby verifying the effectiveness of disentangling.
Note. Additional visualizations are provided in Appendix Secs. 10.2 to 10.4, including more cross-view reconstructions and attention maps.
@@ 4.4 Ablation Study
Effect of CVD’s Components. We conduct ablation studies on the University-1652 dataset to evaluate the contribution of each component in CVD, as summarized in Tab. 5(a). To ensure fair comparison, we adopt two representative pipelines, LPN and Game4Loc, which both use a shared ResNet50 backbone. Removing both constraints (Exp.2 and 7) leads to a notable performance drop, while individually adding the intra-view independence (Exp.3 and 8) or inter-view reconstruction constraint (Exp.4 and 9) yields consistent gains. The best results are obtained when both constraints are jointly applied (Exp.5 and 10), confirming that explicitly factorizing content and viewpoint is essential for robust cross-view alignment.
@@
Analysis of Temperature Parameters. As illustrated in Tab. 5(c), model performance remains stable across a range of temperature values $\tau$. This insensitivity indicates that the performance gains primarily result from effective representation disentanglement rather than contrastive loss tuning.
Effect of Reconstruction Losses. We compare MSE, SSIM, and perceptual losses. As shown in Tab. 6, MSE achieves the highest PSNR/SSIM and best retrieval accuracy, indicating that pixel-level fidelity better preserves view consistency. By contrast, SSIM and perceptual losses relax strict photometric fidelity and show higher tolerance to small misalignments, which may reduce geometric consistency and slightly affect retrieval performance. We therefore adopt MSE in all experiments.
Note. We provide additional ablations in the Appendix, including robustness to training data scale (Tab. 8), the number of SWD projections (Tab. 9), and the effect of loss-balancing weights (Tab. 11), among others.
5 Conclusion
In this paper, we revisit drone-view geo-localization (DVGL) from a manifold-learning perspective and propose CVD, a unified framework that explicitly disentangles content and viewpoint within visual representations. CVD follows an embed-disentangle-reconstruct paradigm, guided by intra-view independence and inter-view reconstruction constraints, to promote content- and viewpoint-specific encoding. Extensive experiments show that CVD consistently improves cross-view matching accuracy across diverse pipelines and enhances robustness and generalization under varied scenarios, viewpoints, and altitudes, while achieving lower inference latency. These results underscore the value of separating content and viewpoint for DVGL and point toward more robust and generalizable DVGL systems. We discuss limitations and directions for future work in the Appendix.